# The $C^{0,1}$ Itô-Ventzell formula for weak Dirichlet processes

### v2

@@ 5 Applications
In this section, we apply Theorem 3.1 to derive novel stochastic calculus formulas for quadratic variations and a representation result for strong solutions of SPDEs.
@@ 5.1 Formulas for quadratic variations
Let $W$ be a one-dimensional Brownian Motion. Then, it holds that
$[|W_{t}|^{3},W_{t}]=3\int_{0}^{t}W_{r}|W_{r}|\mathrm{d}r.$
Proof. First, we define $F(t,x):=x|x|W_{t}$. Then, $F$ is a stochastic flow in the sense of Definition 2.7 with local characteristics $\beta(t,x)=0$ and $\gamma(t,x)=x|x|$. Now, let $K\subset\mathbb{R}$ be compact and $p\geq 1$. Then, we get for any $x,y\in K$:
$\mathbb{E}\left[\sup_{t\in[0,T]}|\gamma(t,x)-\gamma(t,y)|^{p}\right]$ $=\big{|}|x|x-|y|y\big{|}^{p}\leq C(p)|x-y|^{p},$
due to $\gamma$ being locally Lipschitz continuous with (the on a compact set $K$ depending) Lipschitz constant $L_{K}$ and we define $C(p):=L_{K}^{p}$. In particular, Assumption 2 in Theorem 3.1 is fulfilled with this $C(p)$. Since the other assumptions obviously hold, we can apply Theorem 3.1 with $F_{x}(t,x)=2|x|W_{t}$. Then, Theorem 3.1 implies with $X$ being a one-dimensional continuous local martingale:
$X_{t}|X_{t}|W_{t}=0+\int_{0}^{t}X_{r}|X_{r}|\mathrm{d}W_{r}+2\int_{0}^{t}|X_{r}|W_{r}\mathrm{d}X_{r}+\mathcal{B}^{X}(F)_{t},$
where $\mathcal{B}^{X}(F)_{t}$ is a weak zero energy process. Hence, we get for any continuous local martingale $M$:
$[X_{t}|X_{t}|W_{t},M_{t}]=\int_{0}^{t}X_{r}|X_{r}|\mathrm{d}[W,M]_{r}+2\int_{0}^{t}|X_{r}|W_{r}\mathrm{d}[X,M]_{r}.$
Finally, choosing $X:=W$ and $M:=W$, yields the proposition since $\mathrm{d}[W]_{r}=\mathrm{d}r$.
Note that in the previous proof, we could not use the Itô-Ventzell formula in $C^{0,2}$ or $C^{1,2}$ since $F$ is not twice continuously differentiable in the space direction.
Proposition 5.2.
Let $W$ be a one-dimensional Brownian Motion. Then, it holds that
$[((W_{t})_{+})^{2}W_{t},W_{t}]=\int_{0}^{t}((W_{r})_{+})^{2}\mathrm{d}[W]_{r}+2\int_{0}^{t}((W_{r})_{+})W_{r}\mathrm{d}[W]_{r},$
where $x_{+}:=\max\{x,0\}$.
Proof. This proof is analogous to the proof of Proposition 5.1 with $F(t,x):=(x_{+})^{2}W_{t}$ and thus $\beta(t,x)=0$ and $\gamma(t,x)=(x_{+})^{2}$ being again locally Lipschitz.
@@ 5.2 A representation result for strong solutions of SPDEs
In this subsection, the overall target is to give a representation result for strong solutions of stochastic elliptic PDE for a special type of weak Dirichlet processes. Therefore, we start with stating the SPDE giving a generalization of the solution theory (naturally generalizing the theory of Gozzi and Russo [12] who considered PDEs). As in Gozzi and Russo [12], in the elliptical case all of the relating function are independent of $t$ due to the elliptical problem being time-independent.
Let $b:\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $\sigma:\mathbb{R}^{d}\to\mathbb{R}^{d\times d}$ be two continuous functions. Then, we define the elliptic operator $L_{0}:D(L_{0})\subset C^{0}(\mathbb{R}^{d})\to C^{0}(\mathbb{R}^{d})$ with $D(L_{0})=C^{2}(\mathbb{R}^{d})$ by $L_{0}F(x):=b(x)F_{x}(x)+\frac{1}{2}tr(\sigma^{\intercal}(x)F_{xx}(x)\sigma(x)$, where $tr$ denotes the trace of a matrix. In total, we then consider the stochastic inhomogeneous elliptic problem for all $x\in\mathbb{R}^{d}$ and almost all $\omega\in\Omega$:
$L_{0}F(\omega,x)=h(\omega,x),$ (5.1)
where $h(\omega,\cdot)\in C^{0}(\mathbb{R}^{d})$ almost surely.
@@
We say that $F$ with local characteristics $(\beta(x),\gamma(x))$ is a strict solution to the stochastic elliptic problem (5.1) if for a.e. $\omega\in\Omega$ it holds that $F(\omega,\cdot)\in D(L_{0})$, $\beta(\omega,\cdot)\in C^{1}(\mathbb{R})$, $\gamma(\omega,\cdot)\in C^{1+\alpha}(\mathbb{R}^{d})$ for an $\alpha\in(0,1)$, and (5.1) holds.
@@
We say that $F$ is a strong solution to the stochastic elliptic problem 5.1 if there exists a sequence $(F^{n})_{n\in\mathbb{N}}$ with local characteristics $(\beta^{n}(\omega,x),\gamma^{n}(\omega,x))$ and $F^{n}(\omega,\cdot)\in D(L_{0})$ for all $n\in\mathbb{N}$ and a sequence $(h^{n})_{n\in\mathbb{N}}$ with $h^{n}(\omega,\cdot)\in C^{0}(\mathbb{R}^{d})$ for all $n\in\mathbb{N}$ such that for every $n\in\mathbb{N}$ and a.e. $\omega\in\Omega$ holds:
  1. 1.
    $\beta^{n}(\omega,\cdot)\in C^{1}(\mathbb{R})$, and $\gamma^{n}(\omega,\cdot)\in C^{1+\alpha}(\mathbb{R}^{d})$ for an $\alpha\in(0,1)$,
  2. 2.
    $F^{n}$ is a strict solution of $L_{0}F^{n}(\omega,x)=h^{n}(\omega,x)$ for all $x\in\mathbb{R}^{d}$ and a.e. $\omega\in\Omega$,
  3. 3.
    $F^{n}(\omega,\cdot)\xrightarrow{n\to\infty}F(\omega,\cdot)$ and $h^{n}(\omega,\cdot)\xrightarrow{n\to\infty}h(\omega,\cdot)$ in $C^{0}(\mathbb{R}^{d})$ (equipped with the supremum-norm on compact sets), and
  4. 4.
    $\beta^{n}(\omega,\cdot)\xrightarrow{n\to\infty}\beta(\omega,\cdot)$ in $\mathcal{L}^{1}(K;\mathrm{d}t)$ for all compact sets $K\subset\mathbb{R}$.
Note that this a real generalization of the notion of a strict solution since every strict solution is also a strong solution by taking the constant sequences $F^{n}:=F$ and $h^{n}:=h$. The representation theorem states then:
@@
Let $b:\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $\sigma:\mathbb{R}^{d}\to\mathbb{R}^{d\times d}$ be two continuous functions, and $F$ (with local characteristics $(\beta(\omega,x),\gamma(\omega,x))$) be a strong solution of problem (5.1) where $F^{n}$ (with local characteristics $(\beta^{n}(\omega,x),\gamma^{n}(\omega,x))$) denotes the approximating sequence. Define $X_{t}:=x+\int_{0}^{t}\sigma(X_{r})\mathrm{d}r+A_{t}$, where $x\in\mathbb{R}^{d}$ and $A$ is a continuous weak zero energy process such that $X$ is a continuous FQV process. Furthermore, we assume:
  1. 1.
    $F(\omega,\cdot)\in C^{1}(\mathbb{R}^{d})$ for a.e. $\omega\in\Omega$,
  2. 2.
    $\beta$ and $\gamma$ fulfill the assumptions of Theorem 3.1 with $\gamma(\omega,\cdot)\in C^{1}(\mathbb{R}^{d})$,
  3. 3.
@@
    $\lim_{n\to\infty}\int_{0}^{t}(F_{x}(X_{r})-F^{n}_{x}(X_{r}))\mathrm{d}^{-}A_{r}$ $+\int_{0}^{t}(\gamma_{x}(X_{r})-\gamma^{n}_{x}(X_{r}))\sigma(X_{r})\mathrm{d}r$
    $+\int_{0}^{t}(F_{x}(X_{r})-F^{n}_{x}(X_{r}))b(X_{r})\mathrm{d}r=0.$ (5.2)
Then, it holds that (suppressing the dependence on $\omega$)
$F(X_{t})=$ $F(X_{0})+\int_{0}^{t}\gamma(X_{r})\mathrm{d}W_{r}+\int_{0}^{t}F_{x}(X_{r})\sigma(X_{r})\mathrm{d}W_{r}+\int_{0}^{t}\beta(X_{r})\mathrm{d}r+\int_{0}^{t}F_{x}(X_{r})\mathrm{d}^{-}A_{r}$
$+\int_{0}^{t}\gamma_{x}(X_{r})\sigma(X_{r})\mathrm{d}r+\int_{0}^{t}h(X_{r})\mathrm{d}r-\int_{0}^{t}b(X_{r})F_{x}(X_{r})\mathrm{d}r.$
@@
$F^{n}(X_{t})=$ $F^{n}(X_{0})+\int_{0}^{t}\gamma^{n}(X_{r})\mathrm{d}W_{r}+\int_{0}^{t}F^{n}_{x}(X_{r})\sigma(X_{r})\mathrm{d}W_{r}+\int_{0}^{t}\beta^{n}(X_{r})\mathrm{d}r+\int_{0}^{t}F^{n}_{x}(X_{r})\mathrm{d}^{-}A_{r}$
$+\int_{0}^{t}\gamma^{n}_{x}(X_{r})\sigma(X_{r})\mathrm{d}r+\frac{1}{2}\int_{0}^{t}\sigma^{\intercal}(X_{r})F^{n}_{xx}(X_{r})\sigma(X_{r})\mathrm{d}r$
$=$ $F^{n}(X_{0})+\int_{0}^{t}\gamma^{n}(X_{r})\mathrm{d}W_{r}+\int_{0}^{t}F^{n}_{x}(X_{r})\sigma(X_{r})\mathrm{d}W_{r}+\int_{0}^{t}\beta^{n}(X_{r})\mathrm{d}r+\int_{0}^{t}F^{n}_{x}(X_{r})\mathrm{d}^{-}A_{r}$
$+\int_{0}^{t}\gamma^{n}_{x}(X_{r})\sigma(X_{r})\mathrm{d}r+\int_{0}^{t}L_{0}F^{n}(X_{r})\mathrm{d}r-\int_{0}^{t}b(X_{r})F^{n}_{x}(X_{r})\mathrm{d}r.$ (5.3)
@@
$M_{t}^{n}:=$ $F^{n}(X_{t})-F^{n}(X_{0})-\int_{0}^{t}\beta^{n}(X_{r})\mathrm{d}r-\int_{0}^{t}F^{n}_{x}(X_{r})\mathrm{d}^{-}A_{r}-\int_{0}^{t}\gamma^{n}_{x}(X_{r})\sigma(X_{r})\mathrm{d}r$
$-\int_{0}^{t}L_{0}F^{n}(X_{r})\mathrm{d}r+\int_{0}^{t}b(X_{r})F^{n}_{x}(X_{r})\mathrm{d}r,$
$M_{t}:=$ $F(X_{t})-F(X_{0})-\int_{0}^{t}\beta(X_{r})\mathrm{d}r-\int_{0}^{t}F_{x}(X_{r})\mathrm{d}^{-}A_{r}-\int_{0}^{t}\gamma_{x}(X_{r})\sigma(X_{r})\mathrm{d}r$
$-\int_{0}^{t}h(X_{r})\mathrm{d}r+\int_{0}^{t}b(X_{r})F_{x}(X_{r})\mathrm{d}r.$ (5.4)
@@
Note that for almost all $\omega\in\Omega$, it holds that $L_{0}F^{n}(X_{r}(\omega))=h^{n}(X_{r}(\omega))\xrightarrow{n\to\infty}h(X_{r}(\omega))$ uniformly in $r$ due to 5. in Definition 5.4 and $X_{r}(\omega)$ taking values in a compact set for $r\in[0,t]$. Then, by the assumptions of a strong solution and by (3), it holds that $M_{t}^{n}\xrightarrow{n\to\infty}M_{t}$ in $ucp$. Moreover, it follows from (5.3) that $M_{t}^{n}$ is a local martingale. Proposition 4.4 of Gozzi and Russo [12] implies then that $M$ is also a local martingale. Furthermore, note that $M_{t}$ is continuous in $t$ since $X$, $F$, and $A$ are continuous. Then, we define the weak zero energy process
~$\tilde{\mathcal{B}}^{X}(F)_{t}:=$ $\int_{0}^{t}\beta(X_{r})\mathrm{d}r+\int_{0}^{t}F_{x}(X_{r})\mathrm{d}^{-}A_{r}+\int_{0}^{t}\gamma_{x}(X_{r})\sigma(X_{r})\mathrm{d}r+\int_{0}^{t}h(X_{r})\mathrm{d}r-\int_{0}^{t}b(X_{r})F_{x}(X_{r})\mathrm{d}r.$

### v3

@@ 5 Applications
In this section, we apply Theorem 3.1 to derive a novel stochastic calculus formula for quadratic variations and a representation result for strong solutions of SPDEs.
@@ 5.1 Formulas for quadratic variations
In this subsection, we give in Proposition 5.1 a novel stochastic calculus formula for quadratic variations. Afterwards, we point out some special cases. Note that we cannot use the Itô-Ventzell formula in $C^{0,2}$ or $C^{1,2}$ since the function $F$ defined below is not twice continuously differentiable in the space direction. We remark further that Proposition 5.1 is not a special case of the Meyer-Itô formula (see, for instance, Protter [19]) which is only applicable for delta-convex functions, i.e., for the difference of two continuous convex functions. However, the class of $C^{1}$-functions and the class of delta-convex functions are not included in each other, see, e.g., Zelený [27].
@@
Let $W$ be a one-dimensional Brownian Motion, $X=M+A$ be a continuous weak Dirichlet FQV process, and $f:\Omega\times[0,T]\times\mathbb{R}\to\mathbb{R}$ be such that $f(\cdot,x)$ is previsible for all $x\in\mathbb{R}^{d}$ fixed, and continuously differentiable in $x$ for all $t\in[0,T]$. Moreover, we assume that for all $K\subset\mathbb{R}$ compact exists a $C_{K}(\omega)\geq 0$ with $\mathbb{E}[C_{K}]<\infty$ such that $\sup_{(t,x)\in[0,T]\times K}(|f(\omega,t,x)|+|f_{x}(\omega,t,x)|)\leq C_{K}(\omega)$. Then, it holds that
$\left[\int_{0}^{t}f(r,x)\mathrm{d}W_{r}\big{|}_{x=X_{t}},W_{t}\right]=\int_{0}^{t}f(r,X_{r})\mathrm{d}r+\int_{0}^{t}\int_{0}^{r}f_{x}(s,x)\mathrm{d}W_{s}\big{|}_{x=X_{r}}\mathrm{d}[M,W]_{r}.$
Proof. First, we define $F(t,x):=\int_{0}^{t}f(r,x)\mathrm{d}W_{r}$ (where we suppress the $\omega$ as usual). Then, $F$ is a stochastic flow in the sense of Definition 2.7 with local characteristics $\beta(t,x)=0$ and $\gamma(t,x)=f(t,x)$. The assumptions of Theorem 3.1 are fulfilled since continuous differentiability implies local Lipschitz continuity. Moreover, it holds that $F_{x}(t,x)=\int_{0}^{t}f_{x}(r,x)\mathrm{d}W_{r}$ due to the stochastic Leibniz integral rule. Then, Theorem 3.1 implies with the decomposition $X=M+A$:
$\int_{0}^{t}f(r,x)\mathrm{d}W_{r}\big{|}_{x=X_{t}}=0+\int_{0}^{t}f(r,X_{r})\mathrm{d}W_{r}+\int_{0}^{t}\int_{0}^{r}f_{x}(s,x)\mathrm{d}W_{s}\big{|}_{x=X_{r}}\mathrm{d}M_{r}+\mathcal{B}^{X}(F)_{t},$
where $\mathcal{B}^{X}(F)_{t}$ is a weak zero energy process. Hence, we get for any continuous local martingale $\tilde{M}$:
$\left[\int_{0}^{t}f(r,x)\mathrm{d}W_{r}\big{|}_{x=X_{t}},\tilde{M}_{t}\right]=\int_{0}^{t}f(r,X_{r})\mathrm{d}[W,\tilde{M}]_{r}+\int_{0}^{t}\int_{0}^{r}f_{x}(s,x)\mathrm{d}W_{s}\big{|}_{x=X_{r}}\mathrm{d}[M,\tilde{M}]_{r}.$
Finally, choosing $\tilde{M}:=W$ yields the proposition since $\mathrm{d}[W]_{r}=\mathrm{d}r$.
Corollary 5.2.
Let $W$ be a one-dimensional Brownian Motion, and $f:\mathbb{R}\to\mathbb{R}$ be a continuously differentiable function. Then, we get that:
$[f(W_{t})W_{t},W_{t}]=\int_{0}^{t}f(W_{r})\mathrm{d}r+\int_{0}^{t}f^{\prime}(W_{r})W_{r}\mathrm{d}r.$
Proof. This corollary follows immediately from Proposition 5.1 after setting $X=W$ with $\mathrm{d}[W]_{r}=\mathrm{d}r$. Moreover, note that $\int_{0}^{t}\int_{0}^{r}f_{x}(x)\mathrm{d}W_{s}\big{|}_{x=X_{r}}\mathrm{d}[M,W]_{r}=\int_{0}^{t}f^{\prime}(X_{r})W_{r}\mathrm{d}[M,W]_{r}$ in this setting.
For Corollary 5.2, we could have alternatively tried to use an Itô-formula with weaker differentiability conditions.
@@ 5.2 A representation result for strong solutions of SPDEs
In this subsection, the overall target is to give a representation result for strong solutions of stochastic elliptic PDE for a special type of weak Dirichlet processes. We start by giving a generalization of the solution theory of SPDEs (naturally generalizing the theory of Gozzi and Russo [13] who considered PDEs).
Let $b:[0,T]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $\sigma:[0,T]\times\mathbb{R}^{d}\to\mathbb{R}^{d\times d}$ be two continuous functions. Then, we define the time-dependent elliptic operator $L_{0}(t):D(L_{0})\subset C^{0}([0,T]\times\mathbb{R}^{d})\to C^{0}([0,T]\times\mathbb{R}^{d})$ with $D(L_{0})=C^{0,2}([0,T]\times\mathbb{R}^{d})$ by $L_{0}(t)F(t,x):=b(t,x)F_{x}(t,x)+\frac{1}{2}tr(\sigma^{\intercal}(t,x)F_{xx}(t,x)\sigma(t,x)$, where $tr$ denotes the trace of a matrix. We then consider the stochastic inhomogeneous time-dependent elliptic problem for all $(t,x)\in[0,T]\times\mathbb{R}^{d}$ and almost all $\omega\in\Omega$:
$L_{0}(t)F(\omega,t,x)=h(\omega,t,x),$ (5.1)
where $h(\omega,\cdot)\in C^{0}([0,T]\times\mathbb{R}^{d})$ almost surely.
@@
We say that $F$ with local characteristics $(\beta(t,x),\gamma(t,x))$ is a strict solution to the stochastic elliptic problem (5.1) if for a.e. $\omega\in\Omega$ it holds that $F(\omega,\cdot)\in D(L_{0})$, $\beta(\omega,\cdot)\in C^{0,1}([0,T]\times\mathbb{R})$, $\gamma(\omega,\cdot)\in C^{0,1+\alpha}([0,T]\times\mathbb{R}^{d})$ for an $\alpha\in(0,1)$, and (5.1) holds.
@@
We say that $F$ is a strong solution to the stochastic elliptic problem (5.1) if there exists a sequence $(F^{n})_{n\in\mathbb{N}}$ with local characteristics $(\beta^{n}(\omega,t,x),\gamma^{n}(\omega,t,x))$ and a sequence $(h^{n})_{n\in\mathbb{N}}$ such that for every $n\in\mathbb{N}$ and a.e. $\omega\in\Omega$ holds:
  1. 1
    $F^{n}(\omega,\cdot)\in D(L_{0})$, and $h^{n}(\omega,\cdot)\in C^{0}([0,T]\times\mathbb{R}^{d})$,
  2. 2
    $\beta^{n}(\omega,\cdot)\in C^{0,1}([0,T]\times\mathbb{R})$, and $\gamma^{n}(\omega,\cdot)\in C^{0,1+\alpha}([0,T]\times\mathbb{R}^{d})$ for an $\alpha\in(0,1)$,
  3. 3
    $F^{n}$ is a strict solution of $L_{0}(t)F^{n}(\omega,t,x)=h^{n}(\omega,t,x)$ for all $(t,x)\in[0,T]\times\mathbb{R}^{d}$ and a.e. $\omega\in\Omega$,
  4. 4
    $F^{n}(\omega,\cdot)\xrightarrow{n\to\infty}F(\omega,\cdot)$ and $h^{n}(\omega,\cdot)\xrightarrow{n\to\infty}h(\omega,\cdot)$ in $C^{0}([0,T]\times\mathbb{R}^{d})$ (equipped with the supremum-norm on compact sets), and
  5. 5
    $\beta^{n}(\omega,\cdot)\xrightarrow{n\to\infty}\beta(\omega,\cdot)$ in $\mathcal{L}^{1}(K;\mathrm{d}t)$ for all compact sets $K\subset[0,T]\times\mathbb{R}$.
Note that this a generalization of the notion of a strict solution since, of course, every strict solution is also a strong solution by taking the constant sequences $F^{n}:=F$ and $h^{n}:=h$. Our representation theorem states then:
@@
Let $b:[0,T]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $\sigma:[0,T]\times\mathbb{R}^{d}\to\mathbb{R}^{d\times d}$ be two continuous functions, and $F$ (with local characteristics $(\beta(\omega,t,x),\gamma(\omega,t,x))$) be a strong solution of problem (5.1) where we denote the approximation sequence by $F^{n}$ (with local characteristics $(\beta^{n}(\omega,t,x),\gamma^{n}(\omega,t,x))$). Define $X_{t}:=x_{0}+\int_{0}^{t}\sigma(r,X_{r})\mathrm{d}W_{r}+A_{t}$, where $x_{0}\in\mathbb{R}^{d}$ and $A$ is a continuous weak zero energy process such that $X$ is a continuous FQV process. Furthermore, we assume:
  1. 1
    $F(\omega,\cdot)\in C^{0,1}([0,T]\times\mathbb{R}^{d})$ for a.e. $\omega\in\Omega$,
  2. 2
    $\beta$ and $\gamma$ fulfill the assumptions of Theorem 3.1 with $\gamma(\omega,\cdot)\in C^{0,1}([0,T]\times\mathbb{R}^{d})$,
  3. 3
@@
    $\lim_{n\to\infty}\int_{0}^{t}(F_{x}(r,X_{r})-F^{n}_{x}(r,X_{r}))\mathrm{d}^{-}A_{r}$ $+\int_{0}^{t}\mathrm{tr}((\gamma_{x}(r,X_{r})-\gamma^{n}_{x}(r,X_{r}))\sigma(r,X_{r}))\mathrm{d}r$
    $+\int_{0}^{t}(F_{x}(r,X_{r})-F^{n}_{x}(r,X_{r}))b(r,X_{r})\mathrm{d}r=0.$ (5.2)
Then, we have that (suppressing the dependence on $\omega$)
$F(t,X_{t})=$ $F(0,X_{0})+\int_{0}^{t}\gamma(r,X_{r})\mathrm{d}W_{r}+\int_{0}^{t}F_{x}(r,X_{r})\sigma(r,X_{r})\mathrm{d}W_{r}+\int_{0}^{t}\beta(r,X_{r})\mathrm{d}r$
$+\int_{0}^{t}F_{x}(r,X_{r})\mathrm{d}^{-}A_{r}+\int_{0}^{t}\mathrm{tr}(\gamma_{x}(r,X_{r})\sigma(r,X_{r}))\mathrm{d}r+\int_{0}^{t}h(r,X_{r})\mathrm{d}r$
$-\int_{0}^{t}b(r,X_{r})F_{x}(r,X_{r})\mathrm{d}r.$
@@
$F^{n}(t,X_{t})=$ $F^{n}(0,X_{0})+\int_{0}^{t}\gamma^{n}(r,X_{r})\mathrm{d}W_{r}+\int_{0}^{t}F^{n}_{x}(r,X_{r})\sigma(r,X_{r})\mathrm{d}W_{r}+\int_{0}^{t}\beta^{n}(r,X_{r})\mathrm{d}r$
$+\int_{0}^{t}F^{n}_{x}(r,X_{r})\mathrm{d}^{-}A_{r}+\int_{0}^{t}\mathrm{tr}(\gamma^{n}_{x}(r,X_{r})\sigma(r,X_{r}))\mathrm{d}r$
$+\frac{1}{2}\int_{0}^{t}\mathrm{tr}(\sigma^{\intercal}(r,X_{r})F^{n}_{xx}(r,X_{r})\sigma(r,X_{r}))\mathrm{d}r$
$=$ $F^{n}(0,X_{0})+\int_{0}^{t}\gamma^{n}(r,X_{r})\mathrm{d}W_{r}+\int_{0}^{t}F^{n}_{x}(r,X_{r})\sigma(r,X_{r})\mathrm{d}W_{r}+\int_{0}^{t}\beta^{n}(r,X_{r})\mathrm{d}r$
$+\int_{0}^{t}F^{n}_{x}(r,X_{r})\mathrm{d}^{-}A_{r}+\int_{0}^{t}\mathrm{tr}(\gamma^{n}_{x}(r,X_{r})\sigma(r,X_{r}))\mathrm{d}r+\int_{0}^{t}L_{0}(r)F^{n}(r,X_{r})\mathrm{d}r$
$-\int_{0}^{t}b(r,X_{r})F^{n}_{x}(r,X_{r})\mathrm{d}r.$ (5.3)
@@
$M_{t}^{n}:=$ $F^{n}(t,X_{t})-F^{n}(0,X_{0})-\int_{0}^{t}\beta^{n}(r,X_{r})\mathrm{d}r-\int_{0}^{t}F^{n}_{x}(r,X_{r})\mathrm{d}^{-}A_{r}-\int_{0}^{t}\mathrm{tr}(\gamma^{n}_{x}(r,X_{r})\sigma(r,X_{r}))\mathrm{d}r$
$-\int_{0}^{t}L_{0}(r)F^{n}(r,X_{r})\mathrm{d}r+\int_{0}^{t}b(r,X_{r})F^{n}_{x}(r,X_{r})\mathrm{d}r,$
$M_{t}:=$ $F(t,X_{t})-F(0,X_{0})-\int_{0}^{t}\beta(r,X_{r})\mathrm{d}r-\int_{0}^{t}F_{x}(r,X_{r})\mathrm{d}^{-}A_{r}-\int_{0}^{t}\mathrm{tr}(\gamma_{x}(r,X_{r})\sigma(r,X_{r}))\mathrm{d}r$
$-\int_{0}^{t}h(r,X_{r})\mathrm{d}r+\int_{0}^{t}b(r,X_{r})F_{x}(r,X_{r})\mathrm{d}r.$ (5.4)
@@
Note that for almost all $\omega\in\Omega$, it holds that $L_{0}(r)F^{n}(r,X_{r}(\omega))=h^{n}(r,X_{r}(\omega))\xrightarrow{n\to\infty}h(r,X_{r}(\omega))$ uniformly in $r$ due to 4 in Definition 5.4 and $(r,X_{r}(\omega))$ taking values in a compact set for $r\in[0,t]$. Then, by the assumptions of a strong solution and by (3), it holds that $M_{t}^{n}\xrightarrow{n\to\infty}M_{t}$ in $ucp$. Moreover, it follows from (5.3) that $M_{t}^{n}$ is a local martingale. Proposition 4.4 of Gozzi and Russo [13] implies then that $M$ is also a local martingale. Furthermore, note that $M_{t}$ is continuous in $t$ since $X$, $F$, and $A$ are continuous. Then, we define the weak zero energy process
~$\tilde{\mathcal{B}}^{X}(F)_{t}:=$ $\int_{0}^{t}\beta(r,X_{r})\mathrm{d}r+\int_{0}^{t}F_{x}(r,X_{r})\mathrm{d}^{-}A_{r}+\int_{0}^{t}\mathrm{tr}(\gamma_{x}(r,X_{r})\sigma(r,X_{r}))\mathrm{d}r+\int_{0}^{t}h(r,X_{r})\mathrm{d}r$
$-\int_{0}^{t}b(r,X_{r})F_{x}(r,X_{r})\mathrm{d}r.$